Update from Saturday 9th January. By Ian Robins.


D7018 – Graham has now completed fitting the new sealing rubber to the roof panels.

John Cooke and Neil continued electrical testing on the loco and have now successfully operated the
compressor and exhausters.


D6575 – In view of the colder weather forecast, it had been intended to increase the antifreeze
concentration in the cooling system on Saturday but unfortunately the loco had been left in the North
Yard head shunt and due to the work in progress at Williton we were unable to get it moved down the
yard.  See picture.

The last time I tested the coolant it was good down to –8degC so lets hope it doesn’t go past that this
week!

It is hoped it will be moved into the DEPG South Yard during the week.

The loco had been involved with moving a crane wagon around earlier in the week but the Monday
driver had experienced flat batteries necessitating charging.  However, the driver on Tuesday and
Wednesday did not experience any starting difficulties so the situation will be monitored.


D6566 – The lifting gantry was moved into position over the loco inside the shed on Saturday in
preparation for the engine rebuild.

Once the last catch that had been missed was undone, the roof was slid forward towards No.2 cab.

Unfortunately it can only be moved enough to uncover No.1 to 5 cylinders so current thinking is to
install the pistons and heads in those positions and then slide the roof back the other way to do No.6
to 8.

The overhauled components have been laid out next to the loco.

The two cylinder heads acquired from Romania have been denoted with yellow paint to indicate that
they have metric threads, the original British manufactured heads featuring imperial threads.